OpenStudy (anonymous):

What is an example of how you can offer assistance to enhance the overall health of others? Ask your mom to add a vegetable to tonight's dinner to make sure you get a balanced meal. Go with your grandma to her doctor to listen and take notes about her treatment options. Suggest your family hangs out together to watch the new movie that came out on DVD. Tell your brother to get his physical before tennis season starts so he doesn't get cut from the team.

OpenStudy (anonymous):

I think the answer is either A or B because in my notes it says that a way you can make an impact on others is by sharing a well balanced meal and offering to learn about health issues.

OpenStudy (anonymous):

@SolomonZelman @Abhisar @aaronq @beccaboo333 @robtobey @e.mccormick @hba

OpenStudy (anonymous):

@Somy

OpenStudy (solomonzelman):

In D he isn't doing anything so much, because being on the team doesn't by definition mean to improve health (in fact it might mean the opposite) C would be an advice of a psychologist or just a good friend, but not an improvement of a physical health. Despite that it is a good advice mentally. B, I would think that the doctor is taking more action. `Ask your mom to add a vegetable to tonight's dinner to make sure you get a balanced meal.` A Sounds like here he is giving an advice about physical health of his mom I choose A

OpenStudy (solomonzelman):

I am not very good at these topics though.

OpenStudy (anonymous):

ok thanks :) i will let you know if it's correct

OpenStudy (solomonzelman):

I would check before submitting it.

OpenStudy (solomonzelman):

Ask someone who knows more about this.

OpenStudy (anonymous):

well I don't have anyone to ask and your answer was one of the ones i thought it was to begin with

OpenStudy (solomonzelman):

Well, when I give an answer to a problem I usually try to explain it, because when Idk I don't guess but admit it. In this case I had a thought so I gave it .

OpenStudy (abhisar):

I'll choose A because he is doing both sharing a well balanced meal and making her learn about health issues

OpenStudy (abhisar):

I am also inclined towards option A due to the term `overall health`

OpenStudy (anonymous):

A was incorrect, that's okay. :) @Abhisar @SolomonZelman

OpenStudy (abhisar):

Dang ! so it was B ?

OpenStudy (anonymous):

ya. o well. the test was out of 50 questions so its ok

OpenStudy (anonymous):

i got an 85%

OpenStudy (abhisar):

\(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\) Congrats !!

OpenStudy (abhisar):

and sorry for ur loss ! \(\Huge{\overset{\color{#6F4242}{\frown}}{\normalsize \left( \begin{matrix} \color{#005CFF}{•́ \; •̀}\\ \hugeˆ\\ \Large \color{#CD5555}{\smile} \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)